In the following system of equations, what does x equal?
8x + 5y = 5
3x + 3y = 9

Answer by nyc_function(2741)

Solve one of the equations for y and then plug the value you found for y into the other equation to find the value of x.
8x + 5y = 5
5y = -8x + 5
y = (-8x/5) + 5/5
y = (-8x/5) + 1
We plug (-8x/5) + 1 for y in the other equation and solve for x.

3x + 3y = 9
3x + 3(-8x/5) + 1) = 9
3x - (24x/5) + 3 = 9
15x - 24x + 15 = 45
-9x = 45 - 15
-9x = 30
x = 30/-9
x = -30/9
x = -10/3
Done!